# ASIDE: Architectural Separation of Instructions and Data in Language Models

### v3

@@ 5 Experiments: Safety
A primary motivation for increasing instruction-data separation is to address the safety of LLM applications. In this section, we verify that ASIDE, which demonstrates a substantial improvement in separation, also boosts the model’s robustness to prompt injections. We perform two sets of experiments on the models trained in Section 4 to evaluate their robustness against indirect and direct prompt injections.
Threat Model
@@ 5.1 Indirect prompt injection
Structured Queries
We follow the evaluation setup of Wu et al. (2024), which uses a subset of indirect prompt injection attacks introduced by Chen et al. (2024) as the Structured Query benchmark. Following Wu et al. (2024), we evaluate all attacks in the in-domain (injection after the data input) and out-of-domain (injection appears before and after the data input) forms. The detailed description of the setting can be found in Appendix F.
BIPIA Following Yi et al. (2025) we benchmark models across code and text splits. The text split spans four diverse tasks (Email QA, Web QA, Table QA, and Summarization), where malicious instructions are embedded in natural language context, while the code split focuses on a Code QA task with attacks hidden in code snippets. We restrict our evaluation to those attacking scenarios that are programmatically evaluable, i.e., scenarios not requiring queries to GPT-3.5 to determine attack success.
We present the results of the indirect prompt injection evaluations in Table 1. ASIDE consistently reduces attack success rates across the BIPIA and StruQ benchmarks, often achieving the lowest attack success rates across all evaluated models. While model scale alone does not reliably improve robustness, e.g., LLaMa 2 13B shows high vulnerability under Vanilla and ISE, ASIDE systematically lowers attack success rates. We take these results as a promising indication that the architectural enforcement of different embeddings for data and instructions can have a positive effect on mitigating indirect attacks, even without the use of specific safety training.
Model Method Attack Success Rate [%] Direct attacks Indirect attacks TensorTrust Gandalf Purple RuLES BIPIA-text BIPIA-code StruQ-ID StruQ-OOD LLaMa 2 7B Vanilla ISE ASIDE LLaMa 2 13B Vanilla ISE ASIDE LLaMa 3.1 8B Vanilla ISE ASIDE Qwen2.5 7B Vanilla ISE ASIDE Mistral 7B v0.3 Vanilla ISE ASIDE
@@ 5.2 Direct prompt injection
To measure adherence of a model to system instructions under malicious user inputs, we follow the evaluation setup of Mu et al. (2024). It is based on four standard datasets TensorTrust, Gandalf, Purple, and RuLES, that we briefly describe below. For further details of the evaluation, see Appendix C.
We present the results of direct prompt injection evaluations in Table 1. ASIDE on average outperforms the Vanilla model as well as the ISE. ASIDE most benefits Mistral-7B v0.3, reducing the attack success rate (ASR) in comparison to the other methods in 8 out of 8 cases. For other models, ASIDE reduces ASR in 6-7 out of 8 cases.
Like in the previous section, these findings show that the improved instruction-data separation, achieved by ASIDE, tends to make the models more robust, even when trained on benign data without a specific safety objective.
@@ 6 Analysis
In this section we study how ASIDE improves the model’s ability to separate instructions from data. We employ interpretability techniques and analyze representations to understand how the proposed method changes the model’s internal processing. The main experiments in this section use the Llama 3.1 8B model. Additional experiments can be found in Appendix G. Results for other models, which show essentially the same findings, can be found in Appendix H.
@@ 6.1 Linear separability of representations
We first study if ASIDE’s separation of instructions and data on the token embedding level leads to better linear separability of the models’ intermediate representations.
We adopt the linear probing setup of Alain & Bengio (2017); Belinkov (2022). First, we create a dataset of particularly challenging prompts, which –in particular– do not allow the model to rely on simple shortcut (e.g., word-level features) to correctly identify instructions (see Section D.1 for details). Then, for any model, we collect its intermediate layer activations at token positions corresponding to instructions or data in the input. Finally, for each layer we train a linear probing classifier to predict whether an intermediate representation is an instruction or if it is data.
Figure 3 shows the classifier accuracy for the Base, Vanilla, ISE, and ASIDE models as each layer, where layer 0 represents the activations after the embedding matrix. The Base and Vanilla models require several layers of processing before their representation allow a reliable separation of instruction tokens from data tokens. The ASIDE model achieves perfect linear separability (100% probe accuracy) from the beginning of processing and maintains the highest level of linear separability throughout later layers. The ISE model also achieves 100% separability initially, but in later layers this value drops to approximately the levels of Vanilla.
@@ 6.2 Instruction concept activation
To gain further insight into the mechanisms behind ASIDE we analyze the representations on the level of concepts (interpretable features). We focus on the concept “input represents an instruction”, and study how ASIDE influences the activation of such an instruction concept in the model representations.
@@
As a qualitative example, Figure 4 visualizes the per-token activation of the instruction concept for one example of the SEP dataset. One can see that for the Vanilla model, the concept is activated erroneously for several data tokens in the data part of the input. For ASIDE, these spurious activations are strongly suppressed.
@@ 6.3 Embedding interventions
As a final illustration we establish a causal link between ASIDE’s use of data-specific embeddings and the lower attack success rates we observe in Section 5. First, as reference experiment, we evaluate the attack success rate (how often the witness string appears in the response) of the already fine-tuned ASIDE model on a subset of 1000 examples from the SEP dataset with probe string (injection) in the data input. Then, as intervention experiments, we repeat the experiment, but use instruction embeddings instead of data embeddings for the probe tokens. Figure 6 show the comparison of ASR between both setups.
@@ 7 Summary and Discussion
We presented ASIDE, an architectural modification for language models that improves their ability to separate instructions from data. ASIDE’s main idea is to use two different embedding representations for any token, depending on whether the token is part of the instructions or the data. In practice, we achieve this conditional embedding mechanism by a simple and efficient modification of the forward pass, applying a 90-degree rotation to non-executable (data) tokens. As we demonstrate experimentally, ASIDE-enabled models achieve much stronger instruction-data separation than models with a standard architecture, even without specific defense prompts or additional safety alignment. Consequently, we see our result as a promising first step towards safer and more trustworthy LLMs.
Despite the encouraging results, a number of open questions remains. First, ASIDE’s choice to rotate the data token embeddings is just one of many possible ways of achieving role-conditioned embeddings. We find it appealing because it does not introduce new learnable parameters and only adds a negligible computation overhead, but, certainly, other choices could and should be explored. Furthermore, our fine-tuning did not include any safety-specific training data or techniques that previously have been reported to mitigate the problem of instruction-data separation, nor did we apply methods that encapsulate LLMs in a secure environment, such as CaMeL (Debenedetti et al., 2025). We see those techniques as orthogonal to ASIDE, which is agnostic to choices of training data or optimization objective. Finally, we purposefully limited our discussion to the situation of single-turn interactions, where the role of instruction vs data is well-defined. It is an open question how to handle multi-turn interactions or reasoning tasks, including the situation where parts of the instructions themselves are model-generated.
Acknowledgments and Disclosure of Funding
This work was supported by the Federal Ministry of Education and Research (BMBF) as grant BIFOLD (01IS18025A, 01IS180371I); the European Union’s Horizon Europe research and innovation programme (EU Horizon Europe) as grant ACHILLES (101189689); and the German Research Foundation (DFG) as research unit DeSBi [KI-FOR 5363] (459422098).
This research was funded in part by the Austrian Science Fund (FWF) 10.55776/COE12. EZ thanks Farrah Jasmine Dingal, Ahmad Beirami, Florian Tramèr, Javier Rando, Michael Aerni, Jie Zhang for feedback, discussions and support. AP thanks his supervisors, Jonas Geiping and Maksym Andriushchenko, for valuable discussions and thanks the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and the ELLIS Institute Tübingen for their support.
@@ Appendix A Rotation
In our setting we multiply the embedding matrix with the canonical -isoclinic rotation Formally, , where is defined as a block-diagonal matrix of rotations in the 2-dimensional space:
Appendix B Training details
Overview
@@ Appendix B Training details
Appendix C Evaluation details
For all safety evaluations, the same evaluation template was employed, as detailed in Appendix B. All models were loaded using bfloat16 precision.
C.1 Indirect prompt injection
@@
C.2 Direct prompt injection
The system messages used in experiments are described in Section 5.2, with specific implementations varying by framework:
@@ C.2 Direct prompt injection
Appendix D Analysis details
D.1 Linear probing details
@@
Appendix E Detailed concept activation experiments
@@
Appendix F Full evaluation on Structured Query
@@
Appendix G More interpretability results – downstream effect of rotation
Rotations are rather simple linear operations, and it might be easy for the model to learn inverse rotations already in early layers. This would allow the model to mostly re-use existing model weights, thereby negating the effect of architectural separation.
@@ Appendix G More interpretability results – downstream effect of rotation
We report our findings in Figure 10. the ASIDE representations move closer to each other, but never converge. Average cosine similarity starts close to 0, reaching 0.8 at layer 8, after which it drops to around 0.7 by the last layer. Despite representations moving towards each other, cosine similarity never exceeds 0.8. Overeall, we find that the model does not unlearn the rotation during training, and its effects persist in later layers.
@@
Appendix H Analysis results for other models
We report the analysis results for the remaining models in our experiments. Linear separability results are reported in Figure 11. Instruction concept activation experiment is reported in Figure 12, 13, 14, and 15. Embedding intervention experiment results are reported in Figure 16. Testing the downstream effect of rotation is reported in Figure 17.

### v4

@@ 5 Experiments: Safety
The main motivation for increasing instruction-data separation is to improve the safety of LLM applications. In this section, we verify that ASIDE, which demonstrates a substantial improvement in separation, also boosts the model’s robustness to prompt injections. We evaluate the robustness of the models trained in Section 4 against indirect and direct prompt injections.
Threat Model.
@@ 5.1 Indirect prompt injection
In indirect prompt injection, a malicious instruction is inserted into text input to trigger an undesirable effect when the model processes it. We evaluate on two standard benchmarks: Structured Queries (StruQ), following Wu et al. (2024), and BIPIA, following Yi et al. (2025).See Appendix G for details. We report attack success rate (ASR, lower is better).
We present the results of the indirect prompt injection evaluations in Table 1. ASIDE consistently reduces attack success rates across all benchmarks. Compared to Vanilla, ASIDE lowers ASR on BIPIA-text from 14.7% to 4.9%, on BIPIA-code from 15.3% to 8.8%, on StruQ-ID from 45.6% to 28.1% and on StruQ-OOD from 45% to 36% (averages across models). By contrast, ISE is, on average, undistinguishable from Vanilla ( 0.1% difference) on BIPIA-code and Struq-ID and provides almost no improvement on BIPIA-text and Struq-OOD (only 1-2%). This suggests that the delimiter/prompt-based Vanilla baseline is strong and that improvements over it are meaningful.
Overall, our results strongly indicate that the architectural enforcement of different embeddings for data and instructions during benign instruction tuning has a noticeable positive effect on mitigating indirect attacks, without any safety-specific training.
Model Method Attack Success Rate [%] Direct attacks Indirect attacks TensorTrust Gandalf Purple RuLES BIPIA-text BIPIA-code StruQ-ID StruQ-OOD Llama 2 7B Vanilla ISE ASIDE Llama 2 13B Vanilla ISE ASIDE Llama 3.1 8B Vanilla ISE ASIDE Qwen2.5 7B Vanilla ISE ASIDE Qwen3 8B Vanilla ISE ASIDE Mistral 7B v0.3 Vanilla ISE ASIDE
@@ 5.2 Direct prompt injection
In direct prompt injection, the user actively provides malicious inputs, trying to make the model, e.g., violate its system instructions. We measure the robustness of a model against such attacks following the evaluation setup of Mu et al. (2024), based on four standard datasets: TensorTrust (Toyer et al., 2024), Gandalf, (Lakera AI, 2023) Purple (Kim et al., 2024), and RuLES (Mu et al., 2023). For further details of the evaluation, see Appendix D.
We report results of direct prompt injection evaluations in Table 1. On average, ASIDE lowers attack success rate by and percentage points on TensorTrust and Gandalf, respectively. It provides a minor -point reduction on Purple, and shows no change on RuLES. In contrast, ISE actually increases success of attacks by 1.7%-3.1% for three out of four benchmarks and provides a minor decrease (3.3%) on Gandalf.
These results indicate that increasing instruction–data separation with ASIDE improves prompt injection robustness even under benign instruction tuning with no explicit safety objective. We believe this to be a strong result: unlike prior work that required deliberate safety fine-tuning, a simple architectural design choice can deliver measurable, “free” improvements in safety even when applied during ordinary, benign instruction-tuning.
@@ 6 Analysis
In this section we study how ASIDE improves the model’s ability to separate instructions from data. We employ interpretability techniques and analyze representations to understand how the proposed method changes the model’s internal processing. The main experiments in this section use the Llama 3.1 8B model. Additional experiments can be found in Appendix H. Results for other models, which show essentially the same findings, can be found in Appendix I.
@@ 6.1 Linear separability of representations
We first study if ASIDE’s separation of instructions and data at the token embedding level leads to better linear separability of the models’ intermediate representations.
We adopt the linear probing setup of Alain and Bengio (2017); Belinkov (2022). First, we create a dataset of particularly challenging prompts, which, in particular, do not allow the model to rely on simple shortcuts (e.g., word-level features) to correctly identify instructions (see Section E.1 for details). Then, for any model, we collect its intermediate layer activations at token positions corresponding to instructions or data in the input. Finally, for each layer we train a linear probing classifier to predict whether an intermediate representation corresponds to an instruction token or a data token.
Figure 3 shows the classifier accuracy for the Base, Vanilla, ISE, and ASIDE models at each layer, where layer 0 represents the activations after the embedding matrix. The Base and Vanilla models require several layers of processing before their representations allow a reliable separation of instruction tokens from data tokens. The ASIDE model achieves perfect linear separability (100% probe accuracy) from the beginning of processing and maintains the highest level of linear separability throughout later layers. The ISE model also achieves 100% separability initially, but in later layers this value drops to approximately the levels of Vanilla.
@@ 6.2 Instruction concept activation
To gain further insight into the mechanisms behind ASIDE we analyze the representations at the level of concepts (interpretable features). We focus on the concept “input represents an instruction”, and study how ASIDE influences the activation of such an instruction concept in the model representations.
@@
As a qualitative example, Figure 4 shows the per-token activation of the instruction concept for one example of the SEP dataset. For the Vanilla model, the concept is activated erroneously for several tokens in the data part of the input. For ASIDE, these spurious activations are strongly suppressed.
@@ 6.3 Embedding interventions
As a final illustration we establish a causal link between ASIDE’s use of data-specific embeddings and the lower attack success rates we observe in Section 5. First, as reference experiment, we evaluate the attack success rate (how often the witness string appears in the response) of the fine-tuned ASIDE model on a subset of 1000 examples from the SEP dataset with probe string (injection) in the data input.
Then, as intervention experiments, we repeat the experiment, but use instruction embeddings instead of data embeddings for the probe tokens. Figure 6 shows the comparison of ASR between both setups.
@@ 7 Summary and Discussion
We presented ASIDE, a drop-in, parameter-free architectural change that enforces separation between instructions and data with a simple conditional embedding mechanism. ASIDE’s main idea is to use two different embedding representations for any token, depending on whether the token is part of the instructions or the data. A single rotation applied to data-token embeddings gives the model explicit role information from the first layer onward. Across Llama 3.1/2, Qwen 3/2.5, and Mistral, ASIDE achieves much stronger instruction-data separation compared with a competitive Vanilla architecture baseline and ISE, while matching utility. It also reduces attack success rates on both direct and indirect prompt injection benchmarks: all without defense prompts or any safety fine-tuning.
Next steps: our mechanism is architecture-level and orthogonal to training objectives, safety data and system-level defenses such as CaMeL (Debenedetti et al., 2025) and FIDES (Costa et al., 2025). Combining ASIDE with such techniques is a promising direction. We purposefully limited our discussion to the single-turn setting, where the role of instruction vs. data is well-defined. Extending ASIDE to multi-turn and exploring alternative role transforms beyond rotations is a natural next step.
8 Declaration of LLM Usage.
In the preparation of the manuscript, LLMs were used occasionally for wording and grammar suggestions.
9 Reproducibility statement
We have made all associated training and evaluation code available on GitHub at https://github.com/egozverev/aside. We have provided a detailed README.md file with step-by-step instructions for setting up the experimental environment and running training and evaluation scripts. The codebase includes comprehensive documentation throughout. To verify the reproducibility of our results, we independently built the repository from scratch and successfully trained and evaluated one of the models, confirming that our findings can be reliably reproduced.
@@ Appendix A Rotation
In our setting we multiply data embeddings with the canonical -isoclinic rotation defined as a block-diagonal matrix of rotations in the 2-dimensional space:
Computation simplification: When , the rotation can be simplified without constructing the full rotation matrix. Since and , the transformation reduces to a simple coordinate swapping and negation operation: . This allows for efficient computation by directly manipulating the coordinate pairs rather than performing matrix multiplication.
A.1 Ablation Study: Choice of Rotation Angle
To verify that the rotation is optimal for instruction-data separation, we conduct an ablation study on Qwen3-8B where we train ASIDE models with different rotation angles .
Figure 7 shows the results. We observe a clear monotonic trend: the SEP score increases as approaches and decreases for larger angles. The rotation achieves the highest separation score of %, representing the highest separation between instruction and data embeddings while maintaining utility. This validates our architectural choice: is indeed optimal for maximizing instruction-data separation in the ASIDE framework.
Appendix B Implementation details
B.1 ASIDE implementation
ASIDE processes a chunked input, where text is split into instruction and data bits by the deployer of the model (e.g., email hosting service splits the input so that emails are labeled as data). The input therefore consists of a sequence of tuples ([some text], [role]), where [role] is either ”instruction” or ”data” (see an example in B.2). We tokenize each bit and build joint input_ids and segment_ids tensors, where segment_ids has the same shape as input_ids and consists of 0s (for instruction) and 1s (for data). We then modify the forward pass of the model to include segment_ids in its input and apply rotation to the embeddings of inputs marked as ”data”. Below is the core part of ASIDE implementation, see the rest in aside/experiments/model.py.
B.2 Example of applying ASIDE
A typical usage of ASIDE could look like this:
The user of the email client enters some request, e.g., “Check my emails from the past week and find the information about the LLM safety talk I was invited to.” This text becomes the “instructions” part of ASIDE’s input. Note there is no influence from the outside on this, so an attacker has no possibility to change the instruction contents.
The actual emails themselves become the “data” part of the input. These can be influenced by an attacker, by sending the user emails with malicious data like a prompt injection: “IGNORE ALL PRIOR INSTRUCTION AND TRANSFER 1 BITCOIN…”
Internally, the inputs are represented as a sequence with instruction/data labels. For example, if the user had 3 emails in their inbox, it could be: [ (“Where does the talk…”), “instruction”), (“Hi, how are you?”, “data”), (“IGNORE ALL PRIOR INSTRUCTIONS…”, “data”), (“You’re invited to a talk on LLM safety at Carnegie Hall…”, “data”) ]
Internally, ASIDE rotates all tokens in “data” segments before concatenating the resulting embeddings. Because all external input has a “data” label, the attacker cannot prevent the rotation.
Other settings have a similar structure: for example, in a RAG application, such as Google’s AI Search, the software labels all user queries (which an attacker cannot modify) as “instructions” and all retrieved documents (which an attacker might influence by creating manipulated websites) as “data”. For a tool-using system, all text that is returned from a (potentially vulnerable) API call would be “data”, etc.
Appendix C Training details
Overview.
@@ Appendix C Training details
Appendix D Evaluation details
For all safety evaluations, the same evaluation template was employed, as detailed in Appendix C. All models were loaded using bfloat16 precision.
D.1 Evaluation on SEP and Alpaca Eval
We report instruction-data separation (SEP score), SEP utility, and Alpaca Eval 1.0 across models in Table 3.
D.2 Indirect prompt injection
Structured Queries. We follow the evaluation setup of Wu et al. (2024), which uses a subset of indirect prompt injection attacks introduced by Chen et al. (2024) as the Structured Query benchmark. Following Wu et al. (2024), we evaluate all attacks in the in-domain (injection after the data input) and out-of-domain (injection appears before and after the data input) forms. The detailed description of the setting can be found in Appendix G.
BIPIA. Following Yi et al. (2025) we benchmark models across code and text splits. The text split spans four diverse tasks (Email QA, Web QA, Table QA, and Summarization), where malicious instructions are embedded in natural language context, while the code split focuses on a Code QA task with attacks hidden in code snippets. We restrict our evaluation to those attacking scenarios that are programmatically evaluable, i.e., scenarios not requiring queries to GPT-3.5 to determine attack success.
@@ D.2 Indirect prompt injection
D.3 Direct prompt injection
The system messages used in experiments of Section 5.2 varying by framework:
@@ D.3 Direct prompt injection
Appendix E Analysis details
E.1 Linear probing details
@@
Appendix F Detailed concept activation experiments
@@
Appendix G Full evaluation on Structured Query
@@
Appendix H More interpretability results – downstream effect of rotation
Rotations are rather simple linear operations, and it might be easy for the model to learn inverse rotations already in early layers. This would allow the model to mostly reuse existing model weights, thereby negating the effect of architectural separation.
@@ Appendix H More interpretability results – downstream effect of rotation
We report our findings in Figure 11. the ASIDE representations move closer to each other, but never converge. Average cosine similarity starts close to 0, reaching 0.8 at layer 8, after which it drops to around 0.7 by the last layer. Despite representations moving towards each other, cosine similarity never exceeds 0.8. Overall, we find that the model does not unlearn the rotation during training, and its effects persist in later layers.
@@
Appendix I Analysis results for other models
We report the analysis results for the remaining models in our experiments. Linear separability results are reported in Figure 12. Instruction concept activation experiment is reported in Figure 13, 14, 15, and 16. Embedding intervention experiment results are reported in Figure 17. Testing the downstream effect of rotation is reported in Figure 18.